OpenStudy (anonymous):

Point P is located at (−2, 7), and point R is located at (1, 0). Find the y value for the point Q that is located two over three the distance from point P to point R. 4.9 4.7 2.5 2.3

OpenStudy (anonymous):

I think it's 2.3.

OpenStudy (anonymous):

Could someone please check?

OpenStudy (harleygirl18):

P = (-2, 7) and R = (1, 0) A point that is 2/3 of the distance PR from P will be at: -2 + (1-(-2)) * 2/3 , 7 + (0-7) * 2/3 They are asking only for the y-coordinate. We just need to calculate: 7 + (0-7) * 2/3 = 7 - 14/3 = (21 - 14) / 3 = 7/3 = 2.3 so yep your correct

OpenStudy (anonymous):

Thank you!